How would you solve a problem like 4x(x-2)-5x(x-1)=2? When I tried the solution was weird. Could you please help me?
Answer by stanbon(75887)

4x(x-2)-5x(x-1)=2
4x^2-8x-5x^2+5x = 2
-x^2-3x=2
x^2+3x+2 = 0
(x+2)(x+1) = 0
x = -2 or x = -1
====================
Cheers,
Stan H.